# SurfDist: Interpretable Three-Dimensional Instance Segmentation Using Curved Surface Patches

### v1

@@ 4 Discussion
SurfDist learns representations of three-dimensional instances as closed meshes of parameterized curved surface patches. Experimental results demonstrate that it can achieve accuracy superior to that of StarDist-3D with a smaller number of parameters for blob-shaped objects which are common in biomedical imaging data. SurfDist’s experimental success shows that the StarDist architectural approach is effective for training machine learning models to estimate parametric surfaces defining 3D blob-like data, and it motivates further development of computational geometry approaches for three-dimensional instance segmentation.
Despite its state-of-the-art performance on segmentation tasks, SurfDist has two significant limitations which should inform the direction of future work. The first of these is that, like StarDist-3D, SurfDist is ill-suited for datasets where instances are not star-convex, such as branching dendritic cells. Second, additional work is needed to reduce the computational time required for inference using trained SurfDist (or StarDist-3D) models. The bottleneck in inference is the non-maximum suppression step, in which an instance is predicted for every voxel of the input volume and pairwise intersections of predicted instances are computed to de-duplicate predictions. Computation of intersections of objects bounded by polyhedral meshes is a common task in computer graphics and already highly optimized, so a fruitful avenue for reducing inference costs may be development of alternative schemes to non-maximum suppression which decouple instance predictions from input voxel grids.
@@ 5 Data and code availability
Code for the SurfDist model and for reproduction of the experiments detailed here is available at https://github.com/focolab/SurfDist.

### v2

@@ 4 Discussion
SurfDist learns representations of three-dimensional instances as closed meshes of parameterized curved surface patches. Our experimental results demonstrate that it can achieve accuracy superior to that of StarDist-3D with a smaller number of parameters for blob-shaped objects which are common in biomedical imaging data. As the resolution of microscopy increases through technological development, we expect the relative advantage of curved surface modeling over faceted surface models to increase.
Our tests show that the StarDist architectural approach can be effective applied to estimate parametric surfaces underlying 3D blob-like data, and it motivates further incorporation of computational geometry approaches for three-dimensional instance segmentation. For example, SurfDist could be extended to support more complex surface topologies.
Despite its state-of-the-art performance on segmentation tasks, SurfDist has two significant limitations which should inform the direction of future work. First, like StarDist-3D, SurfDist is ill-suited for datasets where instances are not star-convex, such as branching dendritic cells. Second, additional work is needed to reduce the computational time required for inference using trained SurfDist (or StarDist-3D) models. The current bottleneck in inference is the non-maximum suppression step, in which an instance is predicted for every voxel of the input volume and pairwise intersections of predicted instances are computed to de-duplicate predictions. Computation of intersections of objects bounded by polyhedral meshes is a common task in computer graphics and already highly optimized, so a fruitful avenue for reducing inference costs may be development of alternative schemes for non-maximum suppression that decouple instance predictions from input voxel grids.
Finally, our exploration of existing commonly used reference 3D datasets revealed obvious and substantial artifacts in their ground-truth annotations. Limited availability of high-quality training data has long been a serious limiting factor on the advancement of 3D image segmentation approaches, and though efforts toward publication of new datasets have increased encouragingly in recent years, our results indicate that work on dataset generation and curation is still sorely needed. We hope that our results will help to motivate continued work in these areas.
@@ 5 Data and code availability
Code for the SurfDist model and for reproduction of the experiments detailed here is available on GitHub.